OpenStudy (anonymous):

write the standard form of the equation of the line through the pair points (-5,-2) and (-9, 1) Write the equation using only integer coefficients

OpenStudy (bahrom7893):

M=(y2-y1)/(x2-x1) M=(1+2)/(-9+5)=-3/4 y+2=(-3/4)(x+5) Using only integer coefficients: 4y+8=-3x-15 standard form: 4y+3x=-23